Circular No. 8972

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V1309 SCORPII = NOVA SCORPII 2008 S. Nakano, Sumoto, Japan, reports the discovery by K. Nishiyama

(Kurume, Fukuoka-ken) and F. Kabashima (Miyaki-cho, Saga-ken) of a bright star (mag 9.5) on unfiltered CCD images obtained on Sept. 2.4594 UT with a 0.40-m f/9.8 reflector, the presumed nova being located at α = 17^{h}57^{m}32^{s}.93 ± 0^{s}.01, δ = −30^{o}43′10′′.1 ± 0′′.1 (equinox 2000.0). Nothing is visible at this position on their unfiltered CCD frames taken on Aug. 20.476 (limiting mag 12.8) and 21.470 (limiting mag 12.1) using a patrol camera (+ 105-mm-f.l. f/5.6 lens). They note nearby USNO-B1.0-catalogue stars at position end figures 33^{s}.221, 10′′.56 (red mag 12.7) and 33^{s}.015, 10′′.39 (red mag 14.8). Additional independent discoveries also were reported from survey images obtained with digital cameras and telephoto lenses by Y. Sakurai (Mito, Ibaraki-ken, Japan; via Nakano; nova at mag 9.7 on Sept. 3.4) and by Guoyou Sun (Qufu, Shandong, China) and Xing Gao (Urumqi, Xinjiang, China) in the course of the Xingming Observatory Nova Survey (mag ≈ 10.5 on Sept. 2.6). Following posting on the CBAT unconfirmed-objects webpage, several other observers sent observations of V1309 Sco that were included with additional discovery details on CBET 1496. Additional selected magnitudes (mostly from unfiltered CCD images) for the nova: 1958 Apr. 18, [19.0 (red Palomar Sky Survey, via C. Jacques and E. Pimentel, Belo Horizonte, Brazil); 2008 July 30, [12 (Sakurai); Aug. 14, 20, 21, and 30, [13.5: (Sun and Gao); 30.783, [14.5 (D. Chekhovich, S. Korotkiy, and T. Kryachko, Karachay-Cherkessia, Russia); 31, [13.5: (Sun and Gao); Sept. 2.735, 10.5 (Chekhovich et al.); 3.512, 9.0 (Nishiyama and Kabashima); 4.489, 8.3 (Nishiyama and Kabashima); 5.483, 7.1 (Nishiyama and Kabashima). N. N. Samus, Institute of Astronomy, Moscow, informs us that this nova has been assigned the designation V1309 Sco.

H. Naito, Nishi-Harima Astronomical Observatory, reports that lowresolution spectra of V1309 Sco were taken on Sept. 3.49 and 4.47 UT by M. Fujii (Ibara, Okayama, Japan, 0.28-m reflector; range 400–800 nm, resolution R ∼ 500) and on Sept. 5.47 by Naito with the 2.0-m NAYUTA telescope (+ MALLS; range 410–670 nm, R ∼ 1200). The spectra show a smooth continuum with some absorption lines and strong Balmer emission lines, which indicate that this object is indeed a classical nova. The expansion velocities derived from the FWHM of Hα are somewhat slow (670 km/s on Sept. 3.49, 670 km/s on Sept. 4.47, and 470 km/s on Sept. 5.47). The spectrum also shows an interstellar Na D absorption line (EW = 0.6 nm). The ratios of Hα to Hβ are 10 on Sept. 4.47 and 6.7 on Sept. 5.47.

2008 September 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green